# DiscoveryWorld: A Virtual Environment for Developing and Evaluating Automated Scientific Discovery Agents

### v1

@@ 1 Introduction
Our goal is to help remedy these by creating the first virtual discovery environment where solving tasks demands all of the key facets in end-to-end scientific discovery, [^2] While a precise definition of ”scientific discovery” is somewhat elusive [17, 19], we here adopt a pragmatic approach: Our goal is to help develop systems that can perform the end-to-end research process that human scientists engage in as they work on a problem, attempt to answer a question, or more generally advance their field of study. We use the term ”scientific discovery” here to refer to this process. and which covers a broad variety of discovery topics. Our approach is to develop a text-based simulated world (with optional 2D visual overlay), called DiscoveryWorld, where agents can navigate around, interact with objects in the world, use scientific equipment (measuring devices, tools, etc.), and make observations. Agents can then form hypotheses, plan and execute experiments, and draw conclusions to solve challenge tasks developed for this virtual world. DiscoveryWorld tasks are grounded in eight varied topics, such as radioisotope dating, rocket science, and proteomics, to encourage development of agents with general discovery skills rather than hard-wiring to a particular challenge (see Figure 1). The tasks themselves are realistic (but simplified), allowing agents to apply both scientific and commonsense knowledge when attempting them. DiscoveryWorld thus provides an environment for exercising and evaluating general-purpose skills in end-to-end AI discovery systems (see Figure 2).
@@ 3.1 Simulator
[…]
@@ 3.2 Discovery Task Themes, Difficulty Levels, and Parametric Variations
Specific tasks in DiscoveryWorld are parametrically generated from a set of 24 high-level templates. These templates are divided into 8 discovery task themes with 3 levels of difficulty for each theme. A high-level description of the 8 discovery task themes is shown in Table 2, with full details (including spoilers) in Appendix C.3. For each of the 24 {theme $\times$ difficulty} combinations, the DiscoveryWorld simulator is capable of generating a large number of parametric versions of that theme that constitute a particular instance of a task. These parametric variations vary the specific task solution, which typically involves dynamically generating new values for specific object properties, and/or substituting in different task objects. For example, in the Proteomics theme, each parametric variation of a given task generates a different clustering problem, with different data, that points to a different animal as the solution, and places the animals at different starting locations within the environment. Parametric variation generation is deterministic and controlled by a single random seed provided to the task generator. While a large number of specific instantiations of each theme are possible, due to the cost associated with evaluating a large number of tasks, our official benchmark is evaluated on 5 seeds (i.e., random seeds zero through four) for each theme and difficulty level, resulting in a total of $8\times 3\times 5=120$ different game instances or tasks.
@@ D.4 Cost analysis
The full benchmark contains $8$ discovery themes $\times$ 33 difficulty levels $\times$ 55 seeds =120=120 tasks to evaluate. Easy tasks are evaluated to 100 steps, while Normal and Challenge tasks are evaluated to 1000 steps or a hard \$125 limit, whichever came first. This results in a total upper-bound estimate of $40\times 100(easy)+40\times 1000(normal)+40\times 1000(challenge)=84,000$ steps required to complete an evaluation run, assuming no agents finish early due to task completion.
@@ Appendix F Human Scientist Participants
To compare model performance against human performance, we recruited 11 practicing human scientists to complete the DiscoveryWorld tasks, with their performance shown in Table 6. Scientists were recruited on the UpWork platform, each with: (1) an MSc or PhD in a natural science, (2) self-evaluated comfort and fluency with statistical methods and common software like spreadsheets, (3) comfort and previous experience with 2D top-down games. To evaluate the latter, all participants were required to complete a screening task involving a Tutorial in DiscoveryWorld. Participants were given a maximum of one hour to complete each task. While our agent models are run in a zero-shot setting, humans can’t forget their past experiences, so our participants only completed the normal and challenge settings, to prevent any easily-discovered knowledge in the easy setting from providing an advantage in other settings. To collect accurate estimates of task difficulty, all humans completed the same seed (Seed 0) of each task. To evaluate discovered knowledge, participants were asked to keep notes, and explicitly state their hypotheses, and supporting evidence. Due to the varied nature of these notes (some wrote a few words, others entire pages per task) and their modality (text, spreadsheets), discovery knowledge was evaluated manually using the following criteria: $1$ (discovered all discovery knowledge in scorecard), $0.5$ (discovered some discovery knowledge in scorecard), or 0 (did not find critical discovery knowledge). Other metrics (task completion, procedural knowledge) were evaluated automatically. At their option, not all participants completed all tasks.

### v2

@@ 1 Introduction
Our goal is to help remedy these by creating the first virtual discovery environment where solving tasks demands all of the key facets in end-to-end scientific discovery,222 While a precise definition of ”scientific discovery” is somewhat elusive [17, 19], we here adopt a pragmatic approach: Our goal is to help develop systems that can perform the end-to-end research process that human scientists engage in as they work on a problem, attempt to answer a question, or more generally advance their field of study. We use the term ”scientific discovery” here to refer to this process. and which covers a broad variety of discovery topics. Our approach is to develop a text-based simulated world (with optional 2D visual overlay), called DiscoveryWorld, where agents can navigate around, interact with objects in the world, use scientific equipment (measuring devices, tools, etc.), and make observations. Agents can then form hypotheses, plan and execute experiments, and draw conclusions to solve challenge tasks developed for this virtual world. DiscoveryWorld tasks are grounded in eight varied topics, such as radioisotope dating, rocket science, and proteomics, to encourage development of agents with general discovery skills rather than hard-wiring to a particular challenge (see Figure 1). The tasks themselves are realistic (but simplified), allowing agents to apply both scientific and commonsense knowledge when attempting them. DiscoveryWorld thus provides an environment for exercising and evaluating general-purpose skills in end-to-end AI discovery systems (see Figure 2).
@@ 3.1 Simulator
The visual observation contains a screenshot (768px×$\times$512px) similar to the user interface, which provides a $24\times 16$ tile view of the world (each tile is 32px×$\times$32px) centered on the agent (see Fig 4 in the Appendix). This provides information about objects that are farther away than the JSON observation, which is typically limited to a certain distance (configurable, default within 3 tiles) around the agent due to the size (in tokens) of the objects and their associated descriptions.
@@ 3.2 Discovery Task Themes, Difficulty Levels, and Parametric Variations
[…]
@@ D.4 Cost analysis
The full benchmark contains $8$ discovery themes $\times$ $3$ difficulty levels $\times$ $5$ seeds $=120$ tasks to evaluate. Easy tasks are evaluated to 100 steps, while Normal and Challenge tasks are evaluated to 1000 steps or a hard \$125 limit, whichever came first. This results in a total upper-bound estimate of $00040\times 100(easy)+40\times 1000(normal)+40\times 1000(challenge)=84,$ 000 steps required to complete an evaluation run, assuming no agents finish early due to task completion.
@@ Appendix F Human Scientist Participants
To compare model performance against human performance, we recruited 11 practicing human scientists to complete the DiscoveryWorld tasks, with their performance shown in Table 6. Scientists were recruited on the UpWork platform, each with: (1) an MSc or PhD in a natural science, (2) self-evaluated comfort and fluency with statistical methods and common software like spreadsheets, (3) comfort and previous experience with 2D top-down games. To evaluate the latter, all participants were required to complete a screening task involving a Tutorial in DiscoveryWorld. Participants were given a maximum of one hour to complete each task. While our agent models are run in a zero-shot setting, humans can’t forget their past experiences, so our participants only completed the normal and challenge settings, to prevent any easily-discovered knowledge in the easy setting from providing an advantage in other settings. To collect accurate estimates of task difficulty, all humans completed the same seed (Seed 0) of each task. To evaluate discovered knowledge, participants were asked to keep notes, and explicitly state their hypotheses, and supporting evidence. Due to the varied nature of these notes (some wrote a few words, others entire pages per task) and their modality (text, spreadsheets), discovery knowledge was evaluated manually using the following criteria: $1$ (discovered all discovery knowledge in scorecard), $0.5$ (discovered some discovery knowledge in scorecard), or 00 (did not find critical discovery knowledge). Other metrics (task completion, procedural knowledge) were evaluated automatically. At their option, not all participants completed all tasks.